### v1

# Your Paper Title Here

Abstract
We here investigate whether socio-economic indicators, such as household wealth, leave recoverable informational imprints in both satellite imagery (capturing physical features like buildings and roads) and Internet-sourced text (reflecting historical, cultural, and economic narratives of neighborhoods). Using Demographic and Health Survey (DHS) data from African neighborhoods (clusters), we pair high-resolution Landsat images with textual descriptions generated by large language models (LLMs) conditioned on location and year, as well as text retrieved by an LLM-driven AI Search Agent from web sources. We develop a multimodal framework that predicts household wealth (measured by the International Wealth Index (IWI)) through five pipelines: (i) a vision model on satellite images, (ii) an LLM using only location and year, (iii) an AI agent that searches and synthesizes web text, (iv) a joint image-text encoder, and (v) an ensemble of all signals. Our framework yields three contributions. First, evaluations show that fusing vision and agent/LLM-generated text outperforms vision-only baselines in wealth prediction (e.g., $R^{2}=0.77$ vs. 0.630.630.63 on out-of-sample splits), with LLM-internal knowledge (artificial neural memory) proving surprisingly more effective than agent-retrieved text, improving robustness to out-of-country and out-of-time generalization. Second, we find partial representational convergence: fused embeddings from vision and language modalities correlate moderately (median cosine similarity across modalities of about 0.60 after alignment), suggesting a shared latent code of material well-being while retaining complementary details, broadly consistent with the Platonic Representation Hypothesis. Although the superior performance of LLM-only text over agent-retrieved data challenges our Agent-Induced Novelty Hypothesis, modest gains from combining agent data in some splits offer weak support for the idea that agent-gathered information—emerging from dynamic interaction with the Internet—introduces a degree of unique representational structures not fully captured by static LLM knowledge. Third, we release a large-scale multimodal dataset comprising approximately 60,000 DHS clusters, each linked to satellite images, LLM-generated descriptions, and associated texts retrieved by AI agents.
1 Introduction
In an era where the timely and accurate measurement of socio-economic disparities is crucial for effective policy-making, humanitarian aid, and sustainable development, traditional household surveys, such as the Demographic and Health Surveys (DHS), provide invaluable ground-truth data but face significant limitations in terms of scale, cost, and frequency (DHS et al., 2013). These surveys, conducted periodically across low- and middle-income countries, capture wealth metrics such as the International Wealth Index (IWI)—a composite measure of household assets and living conditions—but often leave significant gaps in geographic and temporal coverage, particularly in remote or rapidly changing regions (Sakamoto et al., 2025). To bridge these gaps, researchers have increasingly turned to remote sensing technologies, leveraging high-resolution Earth observation (EO) imagery from satellites, such as Landsat, to infer poverty patterns through visual cues, including infrastructure density, land use, and vegetation health (Yeh et al., 2020; Pettersson et al., 2023; Kakooei et al., 2024a; Burke et al., 2021). While these vision-based approaches have shown promise, they are inherently limited by what can be “seen” from above, often missing nuanced socio-cultural, historical, or contextual factors that influence material well-being (O’Brien, 2023; Zhu et al., 2025).
The advent of large language models (LLMs) and multimodal AI systems offers a complementary pathway, enabling the extraction and synthesis of textual information from vast digital repositories, including web sources and encyclopedic knowledge (Sarmadi et al., 2025). This raises two interrelated questions: (1) To what extent can useful textual information about neighborhoods in low- and middle-income countries be recovered from LLMs’ artificial neural memory or found by AI Search Agents on the Internet? (2) If recoverable, how well can socio-economic status be distilled into a compact, latent representation across vision and language modalities, potentially converging toward a shared “Platonic” representation of material well-being (Huh et al., 2024)? Or, do modalities complement each other to enhance estimation?
Figure 1: DAG representing the imprint of the poverty/wealth index $Y_{i}$ on satellite imagery $\mathbf{M}_{i}$ and textual data $\mathbf{W}_{i}$, processed by LLM’s neural weights and queried by AI Search Agent. $\mathbf{W}_{i}$ is challenging to observe directly, hence colored in red. Instead, the goal is to reconstruct a faithful representation via the LLM’s neural memory or AI agent search capabilities. These texts are denoted $\mathbf{W}_{i}^{\text{Desc}}$ for the LLM, and $\mathbf{W}_{i}^{\text{Trace}}$ (raw search result) and $\mathbf{W}_{i}^{\text{Summary}}$ (summary of search result) for the search agent. $\mathbf{B}_{i}$ represents background factors.
Figure 1 illustrates these questions via a directed acyclic graph (DAG), where the materialization of poverty/wealth index $Y_{i}$ in African neighborhoods causally influences both satellite imagery $\mathbf{M}_{i}$ and textual data $\mathbf{W}_{i}$. Socio-economic conditions leave observable traces in the physical environment (e.g., building density, road networks) visible in EO data, as well as in linguistic artifacts (e.g., historical accounts, news reports) (Daoud et al., 2019; Daoud and Dubhashi, 2023). However, $\mathbf{W}_{i}$ is challenging to observe directly, necessitating LLMs to reconstruct it from neural memory as $\mathbf{W}_{i}^{\text{Desc}}$, and AI Search Agents (ASA) to query the Internet for $\mathbf{W}_{i}^{\text{Trace}}$ (raw traces) and thereafter generate $\mathbf{W}_{i}^{\text{Summary}}$ (summaries).
Despite growing interest in multimodality in poverty mapping, systematic investigations remain scarce, hampered by the lack of aligned, large-scale datasets (Lamichhane et al., 2025; Kakooei et al., 2024b). To address this limitation, we here analyze whether household wealth on the African continent, as measured by DHS IWI scores, can be textually and visually reconstructed and encoded in a joint latent space. Drawing on a continent-scale corpus of more than 60,000 DHS clusters across Africa from 1990 to 2020, we pair high-resolution, cloud-free Landsat composites with LLM-generated spatiotemporal narratives and AI-agent-retrieved contextual information. At its heart, our approach taps five different signals to “see” and describe each neighborhood’s wealth. First, we use satellite images to capture physical clues—things like roads, buildings, and vegetation. Next, we ask a language model to imagine a narrative for the place based only on its location and year. Then, an AI agent goes online, gathers real-world text about the area, and boils it down into a concise summary. After that, we train a joint encoder to blend the visual and textual cues into a single shared representation. Finally, we let an ensemble weigh and combine all five perspectives before we analyze embedding spaces to investigate representational convergence and complementarity across modalities. Figure 2 provides a visual overview.
The remainder of the paper is organized as follows: We first review the related work on remote sensing and AI in poverty mapping. Next, we detail our agent and data curation frameworks, followed by an experimental performance and representational analysis, and conclude with implications, limitations, and future directions.
2 Problem Setup & Related Work
Early efforts in remote sensing for poverty estimation used nighttime light imagery as a proxy for economic activity (Elvidge et al., 2009). More recent advances have focused on daytime satellite imagery combined with deep learning techniques. For example, Jean et al. (2016) demonstrated the use of transfer learning from night lights to predict poverty from daytime images in African countries. Subsequent works have applied convolutional neural networks directly to satellite data for wealth prediction (Yeh et al., 2020; Pettersson et al., 2023; Kakooei et al., 2024b). Interpretability has also been a focus, with methods using object detection to generate explainable poverty maps (Babenko et al., 2017). Recent reviews synthesize the state of Earth observation and ML for poverty research, highlighting applications in causal inference and small-area estimation (Sakamoto et al., 2025). Other studies explore fairness and biases in satellite-based poverty maps (Aiken et al., 2023).
The emergence of large language models (LLMs) has opened new avenues for socio-economic inference using textual data. LLMs have been employed to estimate regional socio-economic indicators directly from prompts (Han et al., 2024). Synergizing LLM agents with knowledge graphs has shown promise for socioeconomic prediction (Zhou et al., 2024). Additionally, biases in LLMs related to socioeconomic attributes have been investigated (Arzaghi et al., 2024; de Pieuchon et al., 2025).
Along with LLMs, multimodal approaches that combine vision and language are increasingly being explored. Sarmadi et al. (2025) leveraged GPT 4’s multimodal capabilities to rank satellite images by poverty levels. Other works combine satellite imagery with non-visual features (such as X/Twitter activity, distance from residential roads, and Internet speed) to improve poverty prediction (Jung et al., 2025). Despite these advances, systematic studies on representational convergence across modalities in poverty mapping remain limited (Lamichhane et al., 2025).
Our work builds on the framework outlined in the Introduction (Figure 1), where the latent poverty/wealth index $Y_{i}$ causally influences both satellite imagery $\mathbf{M}_{i}$ and textual data $\mathbf{W}_{i}$, assuming conditional independence between modalities given $Y_{i}$. This structure motivates our investigation into whether vision and language encoders converge toward a shared “Platonic” representation of material well-being, as inspired by the Platonic Representation Hypothesis as proposed by Huh et al. (2024). While this hypothesis has been examined in general vision-language models (Radford et al., 2021), its application to specific scientific domains, such as economic development and poverty mapping—particularly in contexts where data modalities are gathered endogenously by AI agents—is novel.
Agent-Induced Novelty Hypothesis.
We also introduce in this work a (to our knowledge) new hypothesis—whether dynamic, agent-driven data collection can introduce unique representational structures beyond those captured by static LLM memory. In agentic systems, AI agents autonomously gather and synthesize data via LLM-guided paths, we propose the Agent-Induced Novelty Hypothesis: AI-agent-gathered data injects representational novelty—unique structures emergent from endogenous LLM reasoning and dynamic LLM-data interactions—not fully captured by an implicit assumption in standard Platonic Representation frameworks that data are upstream models.
Indeed, unlike static or LLM-internal representations, agent-induced data evolves through iterative queries, retrievals, and syntheses (Miehling et al., 2025). Does this process introduce complementarities such as real-time contextual adaptations or path-dependent representational shifts? Does this novelty enhance predictive power in tasks like poverty mapping? We will later test the novelty hypothesis by comparing embeddings from agent traces ($\mathbf{W}_{i}^{\text{Trace}}$) against LLM-only descriptions ($\mathbf{W}_{i}^{\text{Desc}}$), evaluating whether fused models exhibit higher complementarity in a way that would indicate novelty separate from Platonic convergence.
Formalization.
To test the Platonic Representation Hypothesis and the Agent-Induced Novelty Hypothesis outlined above, we formalize the poverty mapping task as a supervised regression problem. The objective is to predict the International Wealth Index $Y_{i}\in[0,100]$ for each DHS cluster $i$, which is aggregated at the neighborhood level from household surveys. Let $\mathbf{M}_{i}$ denote the Earth observation (EO) features extracted from satellite imagery (e.g., Landsat multispectral bands, processed via a pre-trained vision model to yield embeddings).
Textual signals, as introduced in the DAG (Figure 1) and elaborated in the hypotheses, are decomposed into: $\mathbf{W}_{i}^{\text{Trace}}$, the raw agent search traces (e.g., concatenated Wikipedia excerpts and web search results); $\mathbf{W}_{i}^{\text{Summary}}$, the agent’s synthesized summary of those traces; $\mathbf{W}_{i}^{\text{Desc}}$, the LLM-only spatiotemporal description (e.g., generated from location-year prompts without external search); and scalar predictions, $\hat{Y}_{i}^{\text{Agent}}$ and $\hat{Y}_{i}^{\text{LLM}}$, from agent and LLM-only decoding, respectively. Embeddings from textual components (e.g., via a language model encoder) are denoted $\mathbf{E}_{i}^{\text{Text}}$, while fused multimodal embeddings are $\mathbf{E}_{i}^{\text{Fused}}=f(\mathbf{M}_{i},\mathbf{E}_{i}^{\text{Text}})$ for some encoding function $f$.
3 Data & Methods
Data Curation.
We create a multimodal dataset centered on DHS units, comprising approximately 60,000 geolocated neighborhood clusters across Africa, with surveys conducted between 1990 and 2020.
For each cluster, we extract the International Wealth Index (IWI) as ground truth, a composite score (0–100) reflecting household assets and conditions, which is aggregated to the cluster (neighborhood) level.
Visual data consist of high-resolution, cloud-free Landsat composites (30m resolution, multispectral bands) centered on DHS coordinates.
Text data are generated via two channels: (i) LLM-generated descriptions (GPT-4.1 Nano, no search or tools enabled) conditioned on year, location coordinates, and location name (reverse-geocoded from coordinates); and (ii) AI-agent-retrieved context, where an LLM-driven agent queries Wikipedia and Internet search to extract socioeconomic, historical, and contextual information about DHS place name, also given year and coordinates.
The AI Search Agent has a GPT-4.1 Nano LLM core, and was constructed using the open-source tool LangGraph (Wang and Duan, 2024); the maximum recursion depth was set to 20, meaning that the agent could take at most 20 steps of iterative searching before completing. In practice, we observed that the majority of agent invocations terminated with three search steps or fewer. (Note that both the LLM and the search agent use the same backbone and identical place data, isolating the impact of agentic capabilities.)
The resulting combined corpus—images, LLM texts, agent traces, with IWI labels, dubbed IWI-Africa-Multimodal—will be released on Hugging Face. [^1] To link with the full DHS data, users must register with the DHS Program and agree to its privacy and data use policies, as outlined at https://dhsprogram.com. See Table 1 for two illustrative AI Search Agent traces.
Table 1: Examples of data gathered by the AI Search Agent.
Lat/Long Place Name Full Agent Trace Wikipedia Trace Search Trace 1/10
-18.85 / 47.58 Manazary, Madagascar Manazary is a small rural commune in the outskirts of Antananarivo... suggesting a moderate level of development typical of rural Madagascar. [concatenated with search results: Antananarivo- Avaradrano is a district... elevation of 1,318 metres... etc.] Manazary is a commune in Madagascar... population estimated at 37,000 in 2001... primarily engaged in agriculture with rice as main crop... 40% in fishing. Antananarivo- Avaradrano is a district of Analamanga in Madagascar. It covers smaller communes in the outskirts of Antananarivo...
-1.63 / 29.36 Kanzenze, Rwanda Rubavu District, including Kanzenze sector, experienced social and economic development... suggests a medium level of wealth and infrastructure. [concatenated with search results: Gender equality on socio-economic development in Rubavu... population and housing census... etc.] Rubavu District is one of seven districts in Western Province, Rwanda... capital is Gisenyi... urban area population of 149,209 in 2012. APPROVAL SHEET This thesis entitled "Gender Equality on socio-economic development in Rubavu district, Rwanda"...
Prediction Framework.
Figure 2 illustrates our implementation of the five pipelines for estimating IWI scores while evaluating convergence and complementarity across modalities.
(i) Vision Pipeline: A Vision model (e.g., a 12-layer Vision Transformer architecture with patch size 16 (Stewart et al., 2022)) is pre-trained unsupervised on Landsat images, then fine-tuned supervised on IWI labels via ridge regression on embeddings.
(ii) LLM-Only Pipeline: An LLM (e.g., Llama-4-Maverick, GPT-4.1 Nano) predicts IWI directly from location-year prompts, leveraging internal neural memory; outputs include prediction, justification of prediction, and confidence in prediction.
(iii) AI Search Agent Pipeline: An AI agent (GPT-4.1 Nano core) retrieves and synthesizes web text (as above), then predicts IWI; we extract traces (raw text) and justifications for embedding.
(iv) Ensemble Pipeline: Pipelines (i)–(iii) independently generate modality-specific embeddings, which are then concatenated and used to train a ridge regression model supervised on IWI labels.
Figure 2: Overview of the quintuple prediction framework for estimating household wealth (IWI) from DHS clusters. The Agent-Induced Novelty Hypothesis questions whether the red arrow is actually present to the same degree as others.
Here, embeddings refer to dense vector representations of input data (e.g., images or text) learned by neural networks, which capture semantic or visual features in a compact, latent space suitable for downstream tasks like regression. In our pipelines, we distinguish between (1) frozen embeddings, which are pre-computed from pre-trained models—e.g., OpenAI’s text-embedding-3-small [abbreviated as OAI] or sentence-transformers/all-mpnet-base-v2, abbreviated as [MPNet]—and not updated during training to preserve general knowledge, and (2) fine-tuned embeddings, which are adjusted on our dataset to adapt to poverty-specific patterns. This distinction allows us to evaluate the trade-offs between leveraging off-the-shelf representations and task-specific optimization.
We also define modality-specific acronyms used across experiments: NMR (Neural Memory Reconstruction) denotes the LLM-only pipeline, where text is generated solely from the model’s internal knowledge (artificial neural memory) based on location and year prompts; ASA (AI Search Agent) refers to the pipeline using agent-retrieved Internet text, including raw traces and summaries; and CV (computer vision) indicates the satellite imagery pipeline.
We conduct three distinct evaluation experiments designed to assess model robustness across spatial and temporal dimensions:
  • Random Split: A standard 80/20 train-test split is applied without any geographic or temporal constraints, serving as the baseline for performance comparison.
  • Out-of-Country (OOC): The model is trained on clusters from a subset of countries (a random 80% subset of the data) and evaluated on held-out countries not present in the training set (the remaining 20%). This test evaluates cross-border generalization and the model’s ability to transfer knowledge beyond national boundaries.
  • Out-of-Time (OOT): The model is trained on data from one time period (encompassing 80% of the data) and evaluated on a disjoint time span (the remaining 20%). This assesses temporal generalization and the model’s resilience to shifts in socio-economic conditions over time.
For each experiment, we ensure strict data partitioning to prevent information leakage. In the OOC split, entire countries are treated as atomic units and assigned to either training or test folds. We randomly assign countries such that the distribution of clusters across folds is balanced. In the OOT split, complete years (e.g., 1990–1995) are used as units of partitioning, ensuring that no overlapping time periods exist between the train and test sets. In the Random Split, cluster assignments are made purely at random, with no restrictions based on location or year.
All models are evaluated using two distinct training strategies, depending on whether embeddings are frozen or fine-tuned:
  • For models with frozen embeddings, we perform 100 bootstrap iterations with an 80/20 train-test split per bootstrap iteration. Final results are reported as the mean and standard error across bootstraps.
  • For models with fine-tuned embeddings, we use 5-fold cross-validation, with a 70/15/15 train/validation/test split per fold. This ensures stable and reliable performance estimates while minimizing overfitting. Uncertainties are estimated as the standard deviation of the performance metrics obtained across the five cross-validation iterations on the test fold.
Each model is evaluated under the same experimental protocol, allowing for fair comparisons between single-modality and multimodal configurations. All evaluations are conducted using single-frame inputs; no temporal image or text sequences are used. This design choice allows us to isolate the impact of modality fusion from temporal dynamics.
Performance is measured using the coefficient of determination ($R^{2}$) and root mean squared error (RMSE) on predicting out-of-sample poverty. The $R^{2}$ metric quantifies the proportion of variance in the International Wealth Index (IWI) that is explained by the model, providing a clear interpretation of the model’s predictive power. Confidence intervals are obtained via bootstrapping (for frozen embedding approaches) or cross-validation (for fine-tuned embedding approaches).
This experimental setup enables an assessment of how multimodal signals—particularly those generated through agent-driven web retrieval and LLM-based reasoning—enhance poverty prediction beyond what is achievable with satellite imagery alone. It also supports our investigation into the representational convergence of vision and language, testing both the Platonic Representation Hypothesis and the Agent-Induced Novelty Hypothesis in a real-world, open-ended context.
4 Results
Figure 3: Test performance ($R^{2}$) across evaluation splits (Random, OOC, OOT), grouped by model procedure and data source. Each dot represents the $R^{2}$ for a model under a specific split strategy. The dot shape encodes the split type, and the dot color encodes the embedding model. Rows are ordered by highest $R^{2}$ for readability. This visualization highlights the performance of various model+embedding combinations across generalization scenarios.
Figure 4: Spatial distribution of the difference in prediction residuals between CV-only and NMR+CV. Blue indicates regions where the baseline CV model outperforms NMR+CV. Red indicates regions where NMR+CV outperforms CV. Hexagons summarize mean residual differences across locations.
General patterns
Our evaluation reveals significant performance gains when combining multi-modal signals for poverty prediction, as shown in Figure 3 (full results in Table A.I.1). The NMR+CV approach using Llama-4-Maverick with OpenAI embeddings achieves the highest performance across all evaluation strategies, with an $R^{2}$ of 0.765 on the random split. This represents a substantial improvement over the best single-modality approach (NMR alone at $R^{2}$ = 0.668) and a significant advancement over the CV-only baseline ($R^{2}$ = 0.634). Our CV-only baseline outperforms the shallow baseline method by Pettersson et al. ($R^{2}$ = 0.60 on OOC split), despite using only daytime satellite imagery (instead of both daytime and nighttime images, which is known to boost performance). This demonstrates the effectiveness of our pipeline and suggests that our multimodal approach can achieve comparable or better results with fewer data sources. The performance gains are consistent across evaluation strategies, with the combined model maintaining strong performance even under OOC and out-of-time (OOT) splits. It is notable that the ASA searchers alone perform below the CV-only benchmark, possibly indicating that finding relevant online information is a noisy process.
Notably, the OOC split (where no countries in the training set appear in the test set) produces the most significant performance degradation across all models, indicating that country-specific features play a critical role in poverty prediction. This is particularly evident in the CV-only baseline, which drops from $R^{2}$ = 0.634 (Random) to $R^{2}$ = 0.446 (OOC). In contrast, the OOT split (where no years in the training set appear in the test set) shows a relatively smaller performance drop, suggesting that year-specific features are less critical than country-specific features for poverty prediction in our dataset.
To further contextualize performance, we conducted an extensive analysis of different model architectures and data sources, building on the textual representations defined earlier: $\mathbf{W}_{i}^{\text{Trace}}$ (raw agent search traces), $\mathbf{W}_{i}^{\text{Summary}}$ (agent-synthesized summaries), and $\mathbf{W}_{i}^{\text{Desc}}$ (LLM-generated descriptions). Here, CleanedTraces refers to all agent-crawled text data ($\mathbf{W}_{i}^{\text{Trace}}$) without filtering; Wikipedia is the subset from Wikipedia sources; JustificationOnly includes only the agent’s justification of IWI prediction text; and JustificationPrediction combines justification with the agent’s scalar prediction. The results reveal that the CleanedTraces approach consistently outperforms other subset text sources, including Wikipedia, JustificationOnly, and JustificationPrediction. This finding suggests that the full breadth of agent-generated context provides richer signals for poverty prediction.
The ASA+CV approach using CleanedTraces achieves an $R^{2}$ of 0.740 on the random split, which is 10.4% higher than the CV-only baseline ($R^{2}$ = 0.634) and 1.8% higher than the best single-modality NMR approach ($R^{2}$ = 0.668). This suggests that the AI-search agent’s raw text collection provides novel information that complements the vision pipeline. See Table A.I.1 for full results.
Our analysis of different embedding models reveals that text-embedding-3-small (OAI) consistently outperforms the fine-tuned MPNet model (sentence-transformers/all-mpnet-base-v2) across all evaluation strategies (see Table A.I.1 and Figure 3 for detailed results). For instance, in the NMR+CV pipeline under random splits, OAI achieves an $R^{2}$ of 0.765 compared to MPNet’s 0.747. MPNet cannot match the performance of the frozen-weights OAI model. This suggests that the pretraining and contextual understanding captured in OAI provide a significant advantage for poverty prediction tasks over a smaller model fine-tuned on our dataset.
Africa-wide Spatial and Temporal Analysis of Model Performance.
To quantify spatial variability in gains from multimodality, Figure 4 plots the following difference in absolute residuals, $(\lvert Y_{i}-\hat{Y}_{i}^{\text{CV}}\rvert-\lvert Y_{i}-\hat{Y_{i}}^{\textrm{Best}}\rvert)$. This value is negative (blue) when the CV baseline outperforms the best multi-modal approach; it is positive (red) otherwise. We benchmark here the best-performing joint model against the computer-vision model, as the CV approach is dominant in the ML literature on poverty prediction.
As visualized in Figure 4, the combined NMR+CV model demonstrates particularly strong performance in densely populated regions such as South Africa, where it achieves improved accuracy compared to the CV-only baseline. It also significantly outperforms in conflict-affected areas of Somalia and central Africa (e.g., Chad). In contrast, performance improvements are more modest along the eastern coast (e.g., Guinea), suggesting that the information provided by the NMR component may be less valuable in regions with moderate levels of development and inequality.
Time-series improvements
To complement this spatial analysis, we also analyzed where the temporal improvements are occurring, using the same metric as for the Africa-wide spatial evaluation. Figure 5 shows that the improvements are consistent over time, where the multi-modal model outperforms the CV-only model. Most of the residual improvements are occurring in the 1990s. Interestingly, this is the period when satellite images are most scarce, with an average pixel availability of about three-quarters across the continent, from 1990 to 2000 Jerzak et al. (2023).
Figure 5: Absolute mean residual difference across years for the best system combining NMR+ASA+CV vs CV-only baseline.
Model Size Analysis.
The performance of different LLMs reveals a clear relationship between model size and performance, with the largest model (Llama-4-Maverick, 405B parameters) achieving the best results. However, the smaller models (GPT-4.1 Nano, Grok-3-Mini) perform comparably well at significantly lower computational cost. This suggests that for practical deployment, moderately sized LLMs can provide an excellent balance between performance and cost.
Agent-Induced Novelty vs. Platonic Representation.
One of the most notable findings from our experiments is that the LLM-only NMR approach consistently outperforms the ASA approaches. For instance, NMR with Llama-4-Maverick achieves $R^{2}$ = 0.668 on Random Split, while the best ASA approach (CleanedTraces) only achieves $R^{2}$ = 0.606, indicating that the LLM’s internal knowledge is more predictive than agent-retrieved text. This contradicts our initial expectation that agent-retrieved data would provide more valuable live context and instead aligns with findings in Gema et al. (2025), which show reduced performance in language tasks as context (e.g., via reasoning chains) becomes overly large.
The marginal gain from NMR+ASA+CV ($R^{2}=0.772$) over NMR+CV in random splits provides modest evidence for the Agent-Induced Novelty Hypothesis, as agent data adds some unique structures not in static LLM knowledge; however, this gain is inconsistent (e.g., absent in OOT splits) and minimal overall, suggesting agent-induced novelty is weak here. These results more strongly support the Platonic Representation Hypothesis, as the LLM’s internal model encodes sufficient shared representations of poverty, with agent data contributing only marginal beneficial novelty in specific scenarios. Future work could quantify this further through embedding distance metrics or ablation studies to isolate agent-specific contributions.
Validation of Intra-Cluster Similarity Against Null Distribution
Figure 6: Histogram of cosine similarities between NMR and CV embeddings, after alignment through canonical correlation analysis (first component), for matched DHS clusters.
To statistically validate the observed alignment in Figure 6, we performed a one-sample $t$-test on the cosine similarities between NMR and CV embeddings for matching clusters, testing the null hypothesis that their mean is zero (with empirical $\sigma\approx 0.11$ derived from random pairings). The test produced a $t$-statistic of 312.96 and a $p$-value $<1\times 10^{-10}$, decisively rejecting the null and confirming significant positive similarities. This evidence supports partial representational convergence across modalities, consistent with the Platonic Representation Hypothesis, as vision and text embeddings share meaningful semantic structures beyond random chance while contributing complementary details to poverty prediction.
Regional Representational Convergence.
Figure 7: Latent representation similarity matrix of matching embeddings from NMR and CV, based on cosine similarity after alignment through canonical correlation analysis (first component). Embeddings are sorted by latitude to allow regional interpretability. See Figure A.I.3 for ASA/CV comparison.
To explore the Platonic Representation Hypothesis in a geographically informed manner, we investigate whether representational alignment between vision/language modalities varies by region, potentially reflecting differences in socio-economic documentation or infrastructure visibility across Africa.
To this end, we compute a similarity matrix between NMR-based and CV-based embeddings using cosine similarity after alignment through canonical correlation analysis (Weenink, 2003), with rows/columns sorted by latitude for regional interpretability (Figure 7). The resulting matrix reveals localized clusters of higher similarity, suggesting that embeddings from geographically proximate regions along similar latitudes exhibit stronger alignment. This spatial coherence indicates that geographically proximate areas exhibit systematically stronger cross-modal alignment.
Additional Robustness Checks.
Appendix I provides details on additional robustness tests. For example, it examines potential training/test leakage. There, in Figure A.I.2, we analyze whether temporal patterns in model performance might indicate leakage into future periods. As shown in Figure A.I.2, $R^{2}$ scores across years for exemplary models (ASA-CleanedTraces-OAI3Small and NMR-Llama4Maverick-OAI3Small) exhibit no clear temporal trend, indicating consistent performance over time and minimal evidence of direct data leakage.
5 Discussion & Conclusion
Overall, our findings support the hypothesis that socio-economic indicators, such as household wealth, can be encoded in compact, recoverable latent representations across visual and textual modalities: fused models outperform single-modality baselines by 21.77% in $R^{2}$ across random and country-holdout evaluations. Analysis within our framework indicates moderate convergence between vision and text embeddings, which exhibit a median cosine similarity of around 0.60 after alignment through canonical correlation analysis—partially aligning with the Platonic Representation Hypothesis (Huh et al., 2024); evidence for agent-induced novelty is limited. By releasing a large-scale multimodal DHS dataset, we enable further research in AI-augmented poverty mapping, with implications for equitable global development monitoring and policy interventions.
From a practical standpoint, the NMR approach offers significant scalability advantages over the ASA approach. The NMR pipeline requires only a single LLM inference per location, whereas the ASA approach requires a full web search and text processing pipeline. This makes NMR more cost-effective and suitable for large-scale poverty mapping applications. Our results show that NMR alone achieves 95% of the performance of the combined NMR+CV approach, making it an excellent choice for resource-constrained deployments.
Limitations of our study include reliance on DHS clusters, which may introduce sampling biases by excluding remote areas from sampling (Kakooei and Daoud, 2024). The search agent may also retrieve post-treatment information from the web, potentially biasing causal analyses. Additionally, the computational demands of agent-driven pipelines at continental scales limit scalability for global or time-series applications.
Future work could integrate causal inference into multimodal socio-economic AI while addressing text-specific challenges (Grimmer et al., 2022; Daoud et al., 2022; Pieuchon et al., 2024)—such as improved detection and control of post-treatment bias in agent representations—to further test the Platonic Representation and Agent-Induced Novelty Hypotheses. $\otimes$